OpenStudy (sherribelle):

I need help with an Geometry activity. I know how to do it but one part I just don't understand. I have to convert the measurements. I'm suppose to create a pool and I have the dimensions. The pool is 22ft = length, 18ft = width, and 90in = height. Now the question is, Find the amount of pool liner material that will be needed to completely line the inside of the pool (sides and bottom). I know exactly how to get the answer, you're suppose to find the Lateral Area, but how would I multiply the feet with inches?

OpenStudy (inkyvoyd):

Convert the inches into feet. use 90/12 feet - otherwise, you could convert the feet into inches, but square feet is a more commonly used number than square inches for such a large surface area.

OpenStudy (sherribelle):

Ok, so it would be 7.5ft. And would the calculations look like this: LA=(80ft)7.5ft LA=600ft^2

OpenStudy (sherribelle):

The reason why I chose 90in for the height is because the next part it says Find the amount of water that will be needed to fill the pool if the water needs to be 6 inches below the top of the pool. 90-6=84 which comes out to 7ft.

OpenStudy (inkyvoyd):

Well, they want you to find the surface area of both the bottom of hte pool and the sides...

OpenStudy (sherribelle):

Minus the top of the pool.

OpenStudy (inkyvoyd):

You'll also want to add in 22*18, but 600 is right for the sides. And yeah, no top of hte pool :P

OpenStudy (sherribelle):

But in order to fine the SA you need to find the LA cause: SA=LA+2B

OpenStudy (reemii):

Indeed inches are useful, because they "talk to you in inches". Since you know that 84inches is 7ft now you can give the volume in ft.

OpenStudy (sherribelle):

Yeah I got the volume, it's: V=2772ft^3 I believe.

OpenStudy (reemii):

yes i think so too.

OpenStudy (inkyvoyd):

That looks right :)

OpenStudy (sherribelle):

So would the amount of sides and bottom put together would be 680ft^2?

OpenStudy (inkyvoyd):

I believe so.

OpenStudy (sherribelle):

Ok, thank you. Now I have to do the dog's pool :/ which is 1/4 the dimensions of the family pool. But I think I got it from here.

OpenStudy (reemii):

(area1 + area2 + area1 +area2) + areabottom. i don't obtain 680ft^2.

OpenStudy (inkyvoyd):

protip: the dimensions of the dog's pool are 1/4 the length, so they are \(\large \frac{1^2}{4^2}=\frac{1}{16}\) the area, and \(\large \frac{1^2}{4^3}=\frac{1}{64}\) the volume of water, *ASSUMING* that the water is filled up to 1.5inches from the top. @reemii, what did you get?

OpenStudy (reemii):

I don't think you can divide the volume of water by 64. for hte dimensions of the pool it's k. I think you have to have a 6in without water at the top. (my area was 996ft^2)

OpenStudy (sherribelle):

Ok thank you, there are 3 steps, and that ^^ was basically step 3 cause it wanted you to show the relationship between the two. For step 2 I have to do the exact same thing as step 1 except with different dimensions.

OpenStudy (inkyvoyd):

Yeah, @reemii is correct about the proportionality for the volume. You'll have to redo it since the dimensions aren't 1/4 (the 6 inches stays the same and doesn't turn into 1.5 inches)

OpenStudy (sherribelle):

Well the height is 90in. minus 6 is 84 which is 7ft. V=Bh B=lw B=22(18) B=396ft^2 V=Bh V=396ft^2(7ft) V=2,772ft^3

OpenStudy (sherribelle):

Yeah, I know that part about the dog's pool cause it says it in the directions. The first part says 1/4 the dimensions of the original, meaning l, w, and h. Then the second part it says find the volume of the pool from bottom to 6in from the top.

OpenStudy (inkyvoyd):

Okay - well your work for the first part looks right to me - if you want to I can check your work for the second part, although to be honest, you seem to know what you are doing :)

OpenStudy (sherribelle):

Thanks, I'm reallllllyyyyy good in math, I just wasn't too sure about the converting part. And I'm working on it now, I'll run it by you in a minute.

OpenStudy (sherribelle):

Ok, so the first part I found the LA and added 20ft. Here's what I did: LA=ph p=2l+2w p=2(5.5)+2(4.5) p=11+9 p=20ft LA=ph LA=20(1.875) LA=37.5ft^2+20ft=57.5ft^2

OpenStudy (sherribelle):

The dimensions for the dog house I multiplied it by 0.25 and got: 5.5ft = length, 4.5ft = width, and 22.5in= height which also equals 1.875ft when you convert it.

OpenStudy (inkyvoyd):

Where'd you get the 20 you added from?

OpenStudy (sherribelle):

It's the perimeter. And you're suppose to find the sides and bottom of the pool just like the last one. So when you find the LA you just found the sides so you add the perimeter which is 20.

OpenStudy (inkyvoyd):

Hmm, you find the LA, but you wanted the area of the bottom which is 4.5*5.5 - why the perimeter?

OpenStudy (sherribelle):

So it would be 24.75ft. What about the last question? The LA is 600ft^2. Instead of adding 80ft (perimeter) you would add 396ft (22*18) which would then equal 996ft^2?

OpenStudy (inkyvoyd):

Uhh, are you talking about the dimensions for the pool? If you are, you had everything right before with 1.875[2(5.5)+2(4.5)], excetp you wanted to add 24.75 to get teh total lining area.

OpenStudy (sherribelle):

Yeah, the whole 996ft^2 thing was for the family pool. But for the dog's pool, the total lining I got 62.25ft^2. But for the volume I got a long number :/

OpenStudy (sherribelle):

Since you're suppose to minus 6 from 22.5 I got 16.5in. So then I tried to find the volume: V=Bh B=lw B=5.5(4.5) B=24.75ft^2 V=Bh V=24.75ft^2(1.375) ----16.5 divide by 12 V=34.03125ft^3

OpenStudy (inkyvoyd):

That looks around right actually.

OpenStudy (inkyvoyd):

I'd round to 34.03, but the number of digits is just cause of dividing resulting in decimals, and multiplying making more decimals.

OpenStudy (sherribelle):

Ohhhh ok lol thank you. Now I have to do step 3

OpenStudy (inkyvoyd):

haha - geometry can be such a pain.

OpenStudy (sherribelle):

Yeah, but it's easier than Algebra. I'm almost done with Geometry, I just have to do the little activities which turn out to take hours :/